What concentration of SO 2 − 3 is in equilibrium with Ag 2 SO 3 ( s ) and 4.80 × 10 − 3 M Ag + ? The K sp of Ag 2 SO 3 can be found in this table.

[ SO 2 − 3 ] = M
Answer #1

ASP Ag, s og = 1,50 x10-14, Ag, so, is) a lAgt (aq) + sor cang), Kop=1.5X15-14 Sparingly soluble salle = 1;5x1014 C Aglam))[SO² (aq)] 15x10-14 (418)²x10-6 = 0,0651 X10-8 / [SO2 (aq)] = 6.51x10tom Concentration of sulphate in soccer) which is in wit